### v1

@@ Abstract
Optical coherence tomography (OCT) is a standard imaging modality in ophthalmology (huang1991optical). Due to its noninvasive nature, it is routinely used to measure alteration in layer thickness and visualize fluid accumulations or other structural changes of the central retina. With this, OCT supports the identification of a broad spectrum of ophthalmological diseases, such as age-related macular degeneration (AMD) (schmidt2017view), central serous chorioretinopathy (CSCR) (chhablani2020multimodal) and diabetic retinopathy (virgili2015optical), as well as neurodegenerative conditions like multiple sclerosis (britze2018optical). Systematic learning-based analysis of OCT, whether of single scans, scan series, or population-based statistics, e.g., based on atlases (chakravarty2018construction), can enhance diagnostics and therapy in clinical practice. However, the nature of its acquisition brings along some specific challenges: In favor of shorter acquisition times, clinical routine OCT uses large inter-B-scan distances, leading to sparse sampling in the slow scanning direction and, consequently, to an anisotropic spatial resolution. Most learning-based approaches using convolutional neural networks (CNNs) therefore opt for two-dimensional (2D) solutions rather than performing full three-dimensional (3D) analyses, e.g., for the segmentation of retinal layers (he2019twoDimOctSeg; pekala2019twoDimOctSeg; li2020deepretina; kepp2023shape). However, 2D segmentations do not take into account the information from neighboring B-scans, which may lead to inconsistencies between segmentations of individual slices. The resulting 3D surfaces typically are irregular and require extensive post-processing using graph-cut approaches (li2006optimal), additional refinement networks (liu2024simultaneous) or advanced regularization techniques (Kepp2019_TopologyPreservingShape). Another problem that arises from the sparsely scanned retina is that small anatomical or pathological structures might be missed, reducing the diagnostic precision of OCT image analysis solutions (velaga2017impact; burchard2024analysis). Volumetric measurements performed on these anisotropic scans are inherently imprecise since structures appear stretched or compressed. Some structural context for the space between the B-scans is maintained by en-face scanning laser ophthalmoscopy (SLO), which offers high-resolution imaging of the fundus and is used as a reference for eye tracking in certain OCT systems. However, SLO does not contain differentiated depth information. Similarly, other 2D modalities, such as color fundus photography or fundus autofluorescence (FAF), can also hold valuable complementary information, especially about pathological structures visible on the retinal surface. The relation between these modalities and visible biomarkers is yet to be fully explored (santarossa2022chronological; velaga2022correlation). The spatial correspondence and the resolution discrepancy between en-face SLO/FAF and OCT B-scans are illustrated in Fig. 1 (left, top and bottom).
@@
The resolution-agnostic nature and structural flexibility of implicit neural representations (INRs) suggest that they have the potential to work well in the presence of these challenges. INRs aim to approximate individual data instances as continuous functions and have gained much traction in research recently. They have been shown to be able to efficiently represent and reconstruct natural images and scenes (sitzmann2020implicit; mildenhall2021nerf). With additional instance-specific latent codes or layer modifications, they become generalizable and can be trained population-based to represent multiple images at once (chen2019learning; park2019deepsdf; kazerouni2024incode; dupont2022data; kim2023generalizable). Consequently, a lot of research has been done to examine their applicability to medical tasks such as semantic segmentation (stolt2023nisf), deformable registration (wolterink2022implicit) or atlas generation (grossbrohmer2024sina; dannecker2024cina). However, these medical applications focus on data with far less anisotropy, such as CT or MRI, while highly anisotropic data like OCT remain rather underexplored.
@@
In this paper, we explore the usage of INRs for OCT-specific tasks focused on the challenging anisotropy of the data (Fig. 1, right). 1) We build upon previous work (kepp2025bridging) and make use of densely sampled but two-dimensional SLO and FAF information for the interpolation in-between B-scans. This increases the resolution in the slow scanning direction and is applied to images from healthy and diseased subjects and called anisotropic interpolation hereafter. We leverage a generalizable INR to incorporate the additional en-face modality and support this task through population-based retinal layer or pathology segmentation. 2) We perform implicit inter-subject registration and combine it with the introduced generalization approach to perform population-based registration and create an intensity- and shape-based atlas. An additional INR network enables the construction of a resolution-independent atlas, making it especially suited for anisotropic image data like OCT.
@@ 2.1 Interpolation of Anisotropic OCT Images
Various interpolation methods exist to enhance anisotropic OCT image data resolution. Simple intensity-based methods are unable to capture shape and localization differences between B-scans, leading to artifacts and ambiguous results. lindberg_2018_octInterpolation propose a weighted combination of linear and transfinite interpolation to achieve an improved interpolation of OCT images, but rely on two OCT images taken with orthogonal scan directions. Registration-based interpolation methods (ehrhardt2007structure) align structural features more accurately and reduce misalignment artifacts, but cannot infer information missing between slices, limiting reconstruction quality in sparsely sampled regions. Generative approaches for super-resolution are capable of producing realistic, detailed images by reconstructing fine textures and complex structures, e.g. (lopez2023SuperResGAN), and thus constitute an approach to B-scan interpolation as well. However, most GAN-based approaches for OCT super-resolution only perform B-scan super-resolution but do not interpolate between them (huang_2019_octSuperRes; das_2020_octSuperRes; yuan_2023_octSuperRes), or require pairs of high- and low-resolution images (huang_2019_octSuperRes; yuan_2023_octSuperRes). Furthermore, generative approaches require a large amount of training data and carry the risk of hallucinations, making them precarious in clinical context (cohen2018distribution).
A limited number of previous works have already used INRs for optical imaging. li2025computational built upon the capability for multi-view scene reconstruction demonstrated by neural radiance fields (NeRF) (mildenhall2021nerf) and proposed a ray tracing-based framework for multi-view reconstruction for OCT of different animal anatomies. Similarly, xiao2025limited used an INR for limited view reconstruction of photoacoustic scans of vessels in mice. These works showed that INRs have the capability to work well with sparsely acquired optical imaging data, and the advances of NeRF can be applied well to the medical field. However, methods for multi-view reconstruction cannot directly be applied to interpolation, since slices have no varied angles or overlaps.
@@ 2.2 INR Super-Resolution
While CNN-based approaches are limited to the resolution of the training data, INRs are resolution-independent, as they deliver continuous representations of the imaged objects. Therefore, INRs inherently are a well-suited choice for the task of super-resolution – which is methodologically close to the here performed B-scan interpolation. Super-resolution of 3D medical images using INRs can, for example, be found in (mcGinnis2023inrSuperRes; wu2022arbitrary; fang2024cycleinr). wu2022arbitrary use a CNN encoder and an INR decoder to generate high-resolution MR images from the respective low-resolution images. Similarly, Fang. et al. use a CNN-based encoder to extract feature vectors from low-resolution CT images. These vectors are fed to an INR together with the 3D coordinates of the high-resolution image to predict the pixel intensities for the higher resolution (fang2024cycleinr). mcGinnis2023inrSuperRes make use of multi-contrast MRI, in which different low-resolution contrasts are acquired in different directions, to produce densely sampled, high-resolution images for single subjects.
While these methods either work with low-resolution but equidistantly sampled images or use more than one anisotropic image of the same scene, the resolution-agnostic nature of INRs also directly allows processing single anisotropic images. However, with very large inter-slice distances, as frequently used in retinal OCT imaging, fitting an INR to a single OCT volume results in overfitting to the positions given in the training data. That is, meaningful results are generated for the B-scan positions used during training, but there is no useful interpolation between the B-scans. In (amiranashvili2022learning), this overfitting problem is overcome with a generalizable INR, which is trained on highly anisotropic segmentations. An MLP is trained that is shared between all training instances, while case-specific latent priors are adapted for each sample, allowing the generation of dense and smooth shapes. Building on this approach, stolt2023nisf train a generalizable INR to reconstruct and segment short-axis cardiac MRI and show smooth segmentations on long-axis images.
@@ 2.3 OCT Atlases
Equivalent to less anisotropic modalities, an OCT atlas can be constructed by combining deformable registration and optimization of an atlas image. In the context of optical imaging, an OCT atlas can be helpful to differentiate between physiological variability in normative tissue and pathological alterations. Based on these differences, many diseases have been shown to be classifiable. chakravarty2018construction created a normative atlas and calculated deviation to classify AMD. They proposed to use a geometric mean image as an initial template and employ discrete registration for computational efficiency due to the large size of OCT B-scans. lee2017atlas employed geometry-based registration to create an atlas and visualized thinning of retinal nerve fiber layers due to glaucoma. khansari2019automated create an atlas based on pathological volumes to measure tissue contraction and expansion due to diabetic retinopathy.
In general, these previous approaches also demonstrated the value of correctly aligning individual layers during the construction of the atlas, sometimes through explicit conditions on layer shapes. However, each of these approaches is limited by the resolution of the original OCT volumes. Consequently, the atlas is highly restrictive regarding the data it is applicable to. With different acquisition setups in clinical practice, this is an important factor to overcome. In this work, we will focus on the methodology to make the atlas generation resolution-agnostic and broadly applicable across varying data settings.
@@ 2.4 INR Registration and Atlases
wolterink2022implicit introduced implicit deformable image registration, where a deformation field is represented by an INR and optimized for a single image pair. There have since been multiple adaptations proposed, e.g., B-spline regularization (sideri2024sinr) or inverse consistency (van2023robust; tian2024nephi). To the best of our knowledge, the applicability of wolterink2022implicit to OCT or similarly anisotropic data has not been explored in the academic literature yet.
INRs can further be used to represent an image atlas through different approaches. grossbrohmer2024sina combine conventional registration with a single INR to represent the atlas image, whereas dannecker2024cina use a generalization approach to represent multiple images with one INR and thus omit any registration.
It is still an open question how well the generalization capability of population-trained INRs translates to medical optical imaging with the associated challenges of sparsely sampled data, small datasets, and low signal-to-noise ratio.
@@ 3.1 Datasets
We use an in-house dataset consisting of a 100 paired OCT volumes and SLO images from left and right eyes of 50 healthy volunteers (one OCT volume and one SLO image per eye), resulting in a total of 100 image pairs, to develop, train, and evaluate generalizable INRs for OCT interpolation, segmentation, and atlas generation. All OCT volumes were acquired with a Spectralis OCT scanner (Heidelberg Engineering) modified for research purposes to allow arbitrary inter-B-scan spacing, but image acquisiton was performed using the angiography acquisition (OCT-A) mode consistent with standard clinical OCT systems. Acquisition in the angiography mode results in slightly reduced B-scan image quality, as fewer repeated scans per position were averaged compared to structural OCT acquired with typical clinical protocols, but provides densely sampled OCT volumes of size voxels covering a field of view of mm3. In Spectralis scanners, an SLO (scanning laser ophthalmoscopy) is used for eye-tracking and acquired routinely alongside the OCT. The SLO images have a size of pixels, covering a larger field of view in the lateral plane (approx. mm2) than the OCT images. To match OCT and SLO, we use the localization information provided by the OCT device to subsample the SLO image at the A-scan positions corresponding to each B-scan. In the OCT images, twelve layers (eleven retinal layers plus choroid) were segmented automatically using the Iowa Reference Algorithms (li2006optimal). Subsequently, segmentation errors were corrected manually by an expert. For preprocessing, all B-scans are flattened in relation to the Bruch’s membrane and cropped to pixels so that the retina is at the center height of the image. Images of right eyes are flipped for identical orientation of all images.
Additional experiments are performed on a clinical dataset containing images from 19 patients with central serous chorioretinopathy (CSCR). The CSCR dataset contains 146 OCT volumes consisting of 25 B-scans with a resolution of pixels. Like the healthy images, the CSCR images span a field of view of mm3. As an additional en-face modality, FAF (fundus autofluorescence) images are used for the CSCR data. They have the same size ( voxels) and field of view ( mm2) as the SLO images. To match the FAF images to the OCT volumes, we use a two-step procedure utilizing the accompanying SLO. First, the FAF image is registered onto the SLO using the MedRegNet registration pipeline (santarossa2022medregnet), with improved RANSAC as proposed in (santarossa2025robust). Next, the registered FAF is subsampled to the OCT resolution using the same SLO localization information as before. This entire two-step procedure, albeit in reverse (i.e., mapping OCT onto FAF), has been detailed in (santarossa2022chronological). For the CSCR data, the retina as well as pathologies (intra- and subretinal fluid and pigment epithelial detachment) were segmented manually by experienced ophthalmologists.
@@ 3.2 Anisotropic Interpolation
To reduce the difficulties arising from the anisotropy of clinical OCT volumes, we build upon the framework for joint inter-B-scan interpolation and segmentation of retinal OCT first proposed in (kepp2025bridging). The method is shown exemplarily for healthy images in Fig. 2. A generalizable INR, extending the methods proposed by amiranashvili2022learning and stolt2023nisf, is used that consists of a multi-layer perceptron (MLP) shared between all training samples and learnable subject-specific priors that condition the MLP on the respective subject. Since small structures between B-scans cannot be represented by the sparsely scanned OCT images alone, we propose to use additional en-face imaging modalities as input to the INR. Here, we exemplarily use SLO and FAF images that are densely scanned in the en-face plane. They therefore provide more detailed information about the localization of structures in the lateral direction than OCT. Although these images have no axial depth, we hypothesize that the lateral information enables the INR to represent structures that are not captured in the OCT scanning protocol and are therefore not visible in the B-scans. The INR with trainable parameters has two separate output layers, one for the reconstruction of the OCT volume and one for the segmentation of anatomical or pathological retinal structures, and is trained with OCT and corresponding en-face images from subjects. The network receives three inputs: a 3D coordinate corresponding to a voxel position in the OCT volume of the subject , the intensity of the en-face image at the respective lateral position , and the subject-specific prior .
@@
to adapt the network parameters and the latent priors simultaneously. Here, is the one-hot encoded ground truth segmentation label at position . A combination of MSE and SSIM loss is used for the reconstruction loss and the binary cross-entropy loss is used for the segmentation loss . As proposed by stolt2023nisf, an L2 regularization loss is applied to penalize large values in the latent codes. The loss components are weighted by .
@@
As shown in our previous work (kepp2025bridging), learning the latent priors for unseen cases this way results in plausible segmentation predictions. The learned latent prior enables both the interpolation of OCT intensities and segmentation labels for any coordinate for which the intensity of the additional 2D modality is known. To do so, a simple forward pass is used to calculate .

### v3

@@ Abstract
Optical coherence tomography (OCT) is a standard imaging modality in ophthalmology (Huang et al., 1991). Due to its noninvasive nature, it is routinely used to measure alteration in layer thickness and visualize fluid accumulations or other structural changes of the central retina. With this, OCT supports the identification of a broad spectrum of ophthalmological diseases, such as age-related macular degeneration (AMD) (Schmidt-Erfurth et al., 2017), central serous chorioretinopathy (CSCR) (Chhablani et al., 2020) and diabetic retinopathy (Virgili et al., 2015), as well as neurodegenerative conditions like multiple sclerosis (Britze and Frederiksen, 2018). Systematic learning-based analysis of OCT, whether of single scans, scan series, or population-based statistics, e.g., based on atlases (Chakravarty et al., 2018), can enhance diagnostics and therapy in clinical practice. However, the nature of its acquisition brings along some specific challenges: In favor of shorter acquisition times, clinical routine OCT uses large inter-B-scan distances, leading to sparse sampling in the slow scanning direction and, consequently, to an anisotropic spatial resolution. Most learning-based approaches using convolutional neural networks (CNNs) therefore opt for two-dimensional (2D) solutions rather than performing full three-dimensional (3D) analyses, e.g., for the segmentation of retinal layers (He et al., 2019; Pekala et al., 2019; Li et al., 2020; Kepp et al., 2023). However, 2D segmentations do not take into account the information from neighboring B-scans, which may lead to inconsistencies between segmentations of individual slices. The resulting 3D surfaces typically are irregular and require extensive post-processing using graph-cut approaches (Li et al., 2006), additional refinement networks (Liu et al., 2024) or advanced regularization techniques (Kepp et al., 2019). Another problem that arises from the sparsely scanned retina is that small anatomical or pathological structures might be missed, reducing the diagnostic precision of OCT image analysis solutions (Velaga et al., 2017; von der Burchard et al., 2024). Volumetric measurements performed on these anisotropic scans are inherently imprecise since structures appear stretched or compressed. Some structural context for the space between the B-scans is maintained by en-face scanning laser ophthalmoscopy (SLO), which offers high-resolution imaging of the fundus and is used as a reference for eye tracking in certain OCT systems. However, SLO does not contain differentiated depth information. Similarly, other 2D modalities, such as color fundus photography or fundus autofluorescence (FAF), can also hold valuable complementary information, especially about pathological structures visible on the retinal surface. The relation between these modalities and visible biomarkers is yet to be fully explored (Santarossa et al., 2022b; Velaga et al., 2022). The spatial correspondence and the resolution discrepancy between en-face SLO/FAF and OCT B-scans are illustrated in Fig. 1 (left, top and bottom).
@@
The resolution-agnostic nature and structural flexibility of implicit neural representations (INRs) suggest that they have the potential to work well in the presence of these challenges. INRs aim to approximate individual data instances as continuous functions and have gained much traction in research recently. They have been shown to be able to efficiently represent and reconstruct natural images and scenes (Sitzmann et al., 2020; Mildenhall et al., 2021). With additional instance-specific latent codes or layer modifications, they become generalizable and can be trained population-based to represent multiple images at once (Chen and Zhang, 2019; Park et al., 2019; Kazerouni et al., 2024; Dupont et al., 2022; Kim et al., 2023). Consequently, a lot of research has been done to examine their applicability to medical tasks such as semantic segmentation (Stolt-Ansó et al., 2023), deformable registration (Wolterink et al., 2022) or atlas generation (Großbröhmer et al., 2024; Dannecker et al., 2024). However, these medical applications focus on data with far less anisotropy, such as CT or MRI, while highly anisotropic data like OCT remain rather underexplored.
@@
In this paper, we explore the usage of INRs for OCT-specific tasks focused on the challenging anisotropy of the data (Fig. 1, right). 1) We build upon previous work (Kepp et al., 2025) and make use of densely sampled but two-dimensional SLO and FAF information for the interpolation in-between B-scans. This increases the resolution in the slow scanning direction and is applied to images from healthy and diseased subjects and called anisotropic interpolation hereafter. We leverage a generalizable INR to incorporate the additional en-face modality and support this task through population-based retinal layer or pathology segmentation. 2) We perform implicit inter-subject registration and combine it with the introduced generalization approach to perform population-based registration and create an intensity- and shape-based atlas. An additional INR network enables the construction of a resolution-independent atlas, making it especially suited for anisotropic image data like OCT.
@@ 2.1 Interpolation of Anisotropic OCT Images
Various interpolation methods exist to enhance anisotropic OCT image data resolution. Simple intensity-based methods are unable to capture shape and localization differences between B-scans, leading to artifacts and ambiguous results. Lindberg et al. (2018) propose a weighted combination of linear and transfinite interpolation to achieve an improved interpolation of OCT images, but rely on two OCT images taken with orthogonal scan directions. Registration-based interpolation methods (Ehrhardt et al., 2007) align structural features more accurately and reduce misalignment artifacts, but cannot infer information missing between slices, limiting reconstruction quality in sparsely sampled regions. Generative approaches for super-resolution are capable of producing realistic, detailed images by reconstructing fine textures and complex structures, e.g. (López-Varela et al., 2023), and thus constitute an approach to B-scan interpolation as well. However, most GAN-based approaches for OCT super-resolution only perform B-scan super-resolution but do not interpolate between them (Huang et al., 2019; Das et al., 2020; Yuan et al., 2023), or require pairs of high- and low-resolution images (Huang et al., 2019; Yuan et al., 2023). Furthermore, generative approaches require a large amount of training data and carry the risk of hallucinations, making them precarious in clinical context (Cohen et al., 2018).
A limited number of previous works have already used INRs for optical imaging. Li et al. (2025) built upon the capability for multi-view scene reconstruction demonstrated by neural radiance fields (NeRF) (Mildenhall et al., 2021) and proposed a ray tracing-based framework for multi-view reconstruction for OCT of different animal anatomies. Similarly, Xiao et al. (2025) used an INR for limited view reconstruction of photoacoustic scans of vessels in mice. These works showed that INRs have the capability to work well with sparsely acquired optical imaging data, and the advances of NeRF can be applied well to the medical field. However, methods for multi-view reconstruction cannot directly be applied to interpolation, since slices have no varied angles or overlaps.
@@ 2.2 INR Super-Resolution
While CNN-based approaches are limited to the resolution of the training data, INRs are resolution-independent, as they deliver continuous representations of the imaged objects. Therefore, INRs inherently are a well-suited choice for the task of super-resolution – which is methodologically close to the here performed B-scan interpolation. Super-resolution of 3D medical images using INRs can, for example, be found in (McGinnis et al., 2023; Wu et al., 2023; Fang et al., 2024). Wu et al. (2023) use a CNN encoder and an INR decoder to generate high-resolution MR images from the respective low-resolution images. Similarly, Fang. et al. use a CNN-based encoder to extract feature vectors from low-resolution CT images. These vectors are fed to an INR together with the 3D coordinates of the high-resolution image to predict the pixel intensities for the higher resolution (Fang et al., 2024). McGinnis et al. (2023) make use of multi-contrast MRI, in which different low-resolution contrasts are acquired in different directions, to produce densely sampled, high-resolution images for single subjects.
While these methods either work with low-resolution but equidistantly sampled images or use more than one anisotropic image of the same scene, the resolution-agnostic nature of INRs also directly allows processing single anisotropic images. However, with very large inter-slice distances, as frequently used in retinal OCT imaging, fitting an INR to a single OCT volume results in overfitting to the positions given in the training data. That is, meaningful results are generated for the B-scan positions used during training, but there is no useful interpolation between the B-scans. In (Amiranashvili et al., 2022), this overfitting problem is overcome with a generalizable INR, which is trained on highly anisotropic segmentations. An MLP is trained that is shared between all training instances, while case-specific latent priors are adapted for each sample, allowing the generation of dense and smooth shapes. Building on this approach, Stolt-Ansó et al. (2023) train a generalizable INR to reconstruct and segment short-axis cardiac MRI and show smooth segmentations on long-axis images.
@@ 2.3 OCT Atlases
Equivalent to less anisotropic modalities, an OCT atlas can be constructed by combining deformable registration and optimization of an atlas image. In the context of optical imaging, an OCT atlas can be helpful to differentiate between physiological variability in normative tissue and pathological alterations. Based on these differences, many diseases have been shown to be classifiable using atlases. Chakravarty et al. (2018) created a normative atlas and calculated deviation to classify AMD. They proposed to use a geometric mean image as an initial template and employ discrete registration for computational efficiency due to the large size of OCT B-scans. Lee et al. (2017) employed geometry-based registration to create an atlas and visualized thinning of retinal nerve fiber layers due to glaucoma. Khansari et al. (2019) create an atlas based on pathological volumes to measure tissue contraction and expansion due to diabetic retinopathy.
In general, these previous approaches demonstrated the value of correctly aligning individual layers during the construction of the atlas, sometimes through explicit conditions on layer shapes. However, each of these approaches is limited by the resolution of the original OCT volumes. Consequently, the atlas is highly restrictive regarding the data it is applicable to. With different acquisition setups in clinical practice, this is an important factor to overcome. In this work, we will focus on the methodology to make the atlas generation resolution-agnostic and broadly applicable across varying data settings.
@@ 2.4 INR Registration and Atlases
Wolterink et al. (2022) introduced implicit deformable image registration, where a deformation field is represented by an INR and optimized for a single image pair. There have since been multiple adaptations proposed, e.g., B-spline regularization (Sideri-Lampretsa et al., 2024) or inverse consistency (Van Harten et al., 2023; Tian et al., 2024). To the best of our knowledge, the applicability of Wolterink et al. (2022) to OCT or similarly anisotropic data has not been explored in the academic literature yet.
INRs can further be used to represent an image atlas through different approaches. Großbröhmer et al. (2024) combine conventional registration with a single INR to represent the atlas image, whereas Dannecker et al. (2024) use a generalization approach to represent multiple images with one INR and thus omit any registration. It is still an open question how well the generalization capability of population-trained INRs translates to medical optical imaging with the associated challenges of sparsely sampled data, small datasets, and low signal-to-noise ratio.
@@ 3.1 Datasets
We use an in-house dataset consisting of paired OCT volumes and SLO images from the left and right eyes of 50 healthy volunteers (one OCT volume and one SLO image per eye), resulting in a total of 100 image pairs, to develop, train, and evaluate generalizable INRs for OCT interpolation, segmentation, and atlas generation. All OCT volumes were acquired with a Spectralis OCT scanner (Heidelberg Engineering) modified for research purposes to allow arbitrary inter-B-scan spacing, but image acquisition was performed using the angiography acquisition (OCT-A) mode consistent with standard clinical OCT systems. Acquisition in the angiography mode results in slightly reduced B-scan image quality, as fewer repeated scans per position are averaged compared to structural OCT acquired with typical clinical protocols, but provides densely sampled OCT volumes of size voxels covering a field of view of mm3. In Spectralis scanners, an SLO (scanning laser ophthalmoscopy) is used for eye-tracking and acquired routinely alongside the OCT. The SLO images have a size of pixels, covering a larger field of view in the lateral plane (approx. mm2) than the OCT images. To match OCT and SLO, we use the localization information provided by the OCT device to subsample the SLO image at the A-scan positions corresponding to each B-scan. In the OCT images, twelve layers (eleven retinal layers plus choroid) were segmented automatically using the Iowa Reference Algorithms (Li et al., 2006). Subsequently, segmentation errors were corrected manually by an expert. For preprocessing, all B-scans are flattened in relation to the Bruch’s membrane and cropped to pixels so that the retina is at the center height of the image. Images of right eyes are flipped for identical orientation of all images.
Additional experiments are performed on a clinical dataset containing images from 19 patients with central serous chorioretinopathy (CSCR). The CSCR dataset contains 146 OCT volumes consisting of 25 B-scans with a resolution of pixels. Like the healthy images, the CSCR images span a field of view of mm3. As an additional en-face modality, FAF (fundus autofluorescence) images are used for the CSCR data. They have the same size ( voxels) and field of view ( mm2) as the SLO images. To match the FAF images to the OCT volumes, we use a two-step procedure utilizing the accompanying SLO. First, the FAF image is registered onto the SLO using the MedRegNet registration pipeline (Santarossa et al., 2022a), with improved RANSAC as proposed in (Santarossa et al., 2025). Next, the registered FAF is subsampled to the OCT resolution using the same SLO localization information as before. This entire two-step procedure, albeit in reverse (i.e., mapping OCT onto FAF), has been detailed in (Santarossa et al., 2022b). For the CSCR data, the retina as well as pathologies (intra- and subretinal fluid and pigment epithelial detachment) were segmented manually by experienced ophthalmologists.
@@ 3.2 Anisotropic Interpolation
To reduce the difficulties arising from the anisotropy of clinical OCT volumes, we build upon the framework for joint inter-B-scan interpolation and segmentation of retinal OCT first proposed in (Kepp et al., 2025). The method is shown exemplarily for healthy images in Fig. 2. A generalizable INR, extending the methods proposed by Amiranashvili et al. (2022) and Stolt-Ansó et al. (2023), is used that consists of a multi-layer perceptron (MLP) shared between all training samples and learnable subject-specific priors that condition the MLP on the respective subject. Since small structures between B-scans cannot be represented by the sparsely scanned OCT images alone, we propose to use additional en-face imaging modalities as input to the INR. Here, we exemplarily use SLO and FAF images that are densely scanned in the en-face plane. They therefore provide more detailed information about the localization of structures in the lateral direction than OCT. Although these images have no axial depth, we hypothesize that the lateral information enables the INR to represent structures that are not captured in the OCT scanning protocol and are therefore not visible in the B-scans. The INR with trainable parameters has two separate output layers, one for the reconstruction of the OCT volume and one for the segmentation of anatomical or pathological retinal structures, and is trained with OCT and corresponding en-face images from subjects. The network receives three inputs: a 3D coordinate corresponding to a voxel position in the OCT volume of the subject , the intensity of the en-face image at the respective lateral position , and the subject-specific prior .
@@
to adapt the network parameters and the latent priors simultaneously. Here, is the one-hot encoded ground truth segmentation label at position . A combination of MSE and SSIM loss is used for the reconstruction loss and the binary cross-entropy loss is used for the segmentation loss . As proposed by Stolt-Ansó et al. (2023), an L2 regularization loss is applied to penalize large values in the latent codes. The loss components are weighted by .
@@
As shown in our previous work (Kepp et al., 2025), learning the latent priors for unseen cases this way results in plausible segmentation predictions. The learned latent prior enables both the interpolation of OCT intensities and segmentation labels for any coordinate for which the intensity of the additional 2D modality is known. To do so, a simple forward pass is used to calculate .